# Clever: A Curated Benchmark for Formally Verified Code Generation

### v3

@@ 4 Related Work
Benchmarks. Recent efforts have developed benchmarks for formal verification with the onset of powerful neural models. FVAPPS [9] uses an LLM on scraped competition problems to automatically create formal specifications for 4715 problems, 1083 of which are guarded with test cases. However, the formal specifications themselves are often easily hackable (see Section A.1), with verification correctness guarded by a layer of test cases. Here, we aim to provide complete formal specifications, which cannot be done accurately with automatic annotation. miniCodeProps [26] contains 201 verification problems regarding data structures and induction problems; however, they do not include specification synthesis or equivalence tests. DafnyBench [27] is a benchmark of 782 stand-alone Dafny programs collected from prior benchmarks and Dafny repositories, where the synthesis task is to generate the verification conditions that allow Dafny to prove correctness. At the time, the best model was Claude 3 Opus which solved $\approx$ 68 % of the problems. Software engineering benchmarks have become extremely popular in recent literature, including benchmarking performance fixing real-world issues [14] and contamination-free code generation [12]. In our work, we employ HumanEval [4] to create Clever, our formal verification and synthesis benchmark. Formal verification is also applied in mathematical domains. Mathlib [28] and the Archive of Formal Proofs [1] constitute formal mathematical repositories in Lean and Isabelle respectively, from which benchmarks have been derived [10, 13]. ProofNet [3] serves as a benchmark for producing proper specifications of mathematical problems. PutnamBench [32] is a formal benchmark of undergraduate-level competition problems in Lean, Isabelle, and Coq.
@@ A.1 FVAPPS Benchmark
The FVAPPS benchmark [9] is another code generation benchmark in Lean. However, unlike Clever, which requires a comprehensive proof of full program behavior, FVAPPS only requires the proof of a limited selection of properties of the program. The limitations of this are illustrated by the FVAPPS example in Figure 7. Here, a problem with a relatively complex natural language description only requires verifying lower-bound and upper-bound properties of the program implementation, as well as a few simple base cases. As can be seen, these properties are provably satisfied by a trivial program that always outputs 0 regardless of the input. Thus, it is clear that only requiring the proof of a small handful of properties does not capture the full intent of the natural language problem. This highlights the necessity of a verified code generation benchmark to require proofs of full program behavior, not just program properties.
{mdframed}
[roundcorner=10pt]
solve_elections:
There are n voters, and two ways to convince each of them to vote for you. The first way to convince the \$i\$-th voter is to pay him \$p_i\$ coins. The second way is to make \$m_i\$ other voters vote for you, and the \$i\$-th voter will vote for free. Moreover, the process of such voting takes place in several steps. For example, if there are five voters with \$m_1 = 1\$, \$m_2 = 2\$, \$m_3 = 2\$, \$m_4 = 4\$, \$m_5 = 5\$, then you can buy the vote of the fifth voter, and eventually everyone will vote for you. Set of people voting for you will change as follows: \$5\$ \$1, 5\$ \$1, 2, 3, 5\$ \$1, 2, 3, 4, 5\$. Calculate the minimum number of coins you have to spend so that everyone votes for you.
def solve_elections (n : Nat) (voters : List (Nat × Nat)) : Nat := 0
theorem solve_elections_nonnegative (n : Nat) (voters : List (Nat × Nat)) : solve_elections n voters >= 0 :=
by rfl
theorem solve_elections_upper_bound (n : Nat) (voters : List (Nat × Nat)) : solve_elections n voters <= List.foldl acc (pair : Nat × Nat) => acc + pair.2) 0 voters :=
Nat.zero_le _
theorem solve_elections_zero_votes (n : Nat) (voters : List (Nat × Nat)) : (List.all voters (fun pair => pair.1 = 0)) -> solve_elections n voters = 0 :=
fun _ => rfl
theorem solve_elections_single_zero_vote : solve_elections 1 [(0, 5)] = 0 :=
by rfl
@@ A.2 Hard to write Specifications
Figure 8 shows some problems for which the formal specification or the implementation is hard to write.
{mdframed}
[roundcorner=10pt] (a) def problem_spec -- function signature (implementation: List Rat Rat) -- inputs (xs: List Rat) := -- spec let spec (result: Rat) := let eps := (1: Rat) / 1000000; xs.length 1 xs.length % 2 = 0 poly : Polynomial Rat, poly.degree = some (xs.length - 1) (∀ i, i xs.length - 1 poly.coeff i = xs.get! i) |poly.eval result| eps; -- program termination result, implementation xs = result spec result -- possible implementation using Newtons method def implementation (xs: List Rat) : Rat := let rec poly (xs: List Rat) (x: Rat) := xs.reverse.foldl acc a => acc * x + a) 0; let rec poly (xs: List Rat) (x: Rat) := (xs.drop 1).reverse.foldl acc a => acc * x + a) 0; let rec eps := (1: Rat) / 1000000; let rec find_zero (xs: List Rat) (guess: Rat) (fuel: Nat) := let eval := poly xs guess; let eval := poly xs guess; if eval eps fuel = 0 then (guess, fuel) else let guess := (eval * guess - eval) / eval’; find_zero xs guess (fuel - 1); (find_zero xs 1.0 1000000).1 (b) def problem_spec -- function signature (implementation: Nat Nat) -- inputs (n: Nat) := -- spec let spec (result: Nat) := n > 0 (∃ i, Nat.fib i = result Nat.Prime result (∃! S : Finset Nat, S.card = n - 1 (∀ y S, (∃ k, y = Nat.fib k) y < result Nat.Prime y)) ) -- implementation without proof of -- termination def implementation (n: Nat) : Nat := let rec fib_prime (n: Nat) (i: Nat) : Nat := if Nat.Prime (Nat.fib i) then if n = 1 n = 0 then Nat.fib i else fib_prime (n - 1) (i + 1) else fib_prime n (i + 1) termination_by n decreasing_by -- Proof of termination is open problem sorry sorry fib_prime n 0
@@ A.3 Writing non-computable specifications
Figure 9 shows a computable vs non-computable version of the specification for finding the $n^{th}$ Fibonacci number. It can be observed that the computable version of the specification leaks the implementation in contrast to the non-computable version. The non-computable specification uses an inductive definition of a recursive function.
{mdframed}
[roundcorner=10pt] (a) -- computable spec def problem_spec -- function signature (implementation: List Nat Nat) -- inputs (n: Nat) := -- spec let spec (result: Nat) := (n = 0 result = 0) (n = 1 result = 1) (2 n fib_array : List Nat, fib_array.length = n + 1 fib_array[0]! = 0 fib_array[1]! = 1 (∀ i, 1 < i i < n + 1 fib_array[i]! = fib_array[i - 1]! + fib_array[i - 2]!) result = fib_array[n]!) -- program termination result, implementation xs = result spec result (b) -- non-computable spec inductive fibonacci_non_computable : Prop | base0 : fibonacci_non_computable 0 0 | base1 : fibonacci_non_computable 1 1 | step : n f f₂, fibonacci_non_computable n f fibonacci_non_computable (n + 1) f fibonacci_non_computable (n + 2) (f + f₂) def problem_spec -- function signature (implementation: Nat Nat) -- inputs (n: Nat) := -- spec let spec (result: Nat) := fibonacci_non_computable n result -- program termination result, implementation xs = result spec result
@@
Writing non-computable specifications is a non-trivial task that requires a deep understanding of the problem. Figure 10 (problem 160) presents another complex example illustrating the difficulty of formulating such specifications. Figure 10 shows two versions of a specification for evaluating an expression given as a list of strings (["2","+","3","*","4","-","5"]). Figure 10(a) evaluates the expression and later checks if the output matches the result (not specified in the figure), which is computable. Figure 10(b) shows a non-computable version of the specification that checks if the result matches the output of evaluating the expression without leaking the implementation. One can notice that we need multiple inductive recursive definitions to ensure that the specification is clean and non-computable.
{mdframed}
@@
[roundcorner=10pt] (a) inductive Op where | add | sub | mul | floordiv deriving Repr, DecidableEq def parseOp : String Option Op | "+" => some .add | "-" => some .sub | "*" => some .mul | "//" => some .floordiv | _ => none def precedence : Op Nat | .mul | .floordiv => 2 | .add | .sub => 1 def apply : Op Int Int Int | .add, a, b => a + b | .sub, a, b => a - b | .mul, a, b => a * b | .floordiv, a, b => a / b inductive Tok where | num : Int Tok | op : Op Tok deriving Repr def tokenize : List String Option (List Tok) | [] => some [] | s :: t => match parseOp s with | some o => (tokenize t).map (Tok.op o :: ·) | none => s.toInt?.bind (fun n => (tokenize t).map (Tok.num n :: ·)) partial def evalPass (xs : List Tok) (ops : List Op) : List Tok := match xs with | Tok.num a :: Tok.op o :: Tok.num b :: rest => if o ops then evalPass (Tok.num (apply o a b) :: rest) ops else Tok.num a :: Tok.op o :: evalPass (Tok.num b :: rest) ops | x :: xs => x :: evalPass xs ops | [] => [] def evalTokens (tokens : List Tok) : Option Int := let result := [[.mul, .floordiv], [.add, .sub]].foldl evalPass tokens match result with | [Tok.num n] => some n | _ => none def do_algebra (input : List String) : Option Int := tokenize input >>= evalTokens (b) def applyOp (x y : Int) : String Option Int | "+" => some (x + y) | "-" => some (x - y) | "*" => some (x * y) | "//" => if y == 0 then none else some (x / y) | _ => none inductive evalArith_pass : List String Int Prop | num {s : String} {n : Nat} (h : s.toNat! = n) : evalArith_pass [s] (Int.ofNat n) | binOp {ts1 ts2 : List String} {op : String} {r1 r2 r : Int} (h1 : evalArith_pass ts1 r1) (h2 : evalArith_pass ts2 r2) (hop : applyOp r1 r2 op = some r) : evalArith_pass (ts1 ++ op :: ts2) r inductive evalArith_mul : List String Int Prop | of_pass {ts r} (h : evalArith_pass ts r) : evalArith_mul ts r | step {ts1 ts2 r1 r2 r} (h1 : evalArith_mul ts1 r1) (h2 : evalArith_mul ts2 r2) (hop : applyOp r1 r2 "*" = some r applyOp r1 r2 "//" = some r) : evalArith_mul (ts1 ++ "*" :: ts2) r inductive evalArith_add : List String Int Prop | of_mul {ts r} (h : evalArith_mul ts r) : evalArith_add ts r | step {ts1 ts2 r1 r2 r} (h1 : evalArith_add ts1 r1) (h2 : evalArith_add ts2 r2) (hop : applyOp r1 r2 "+" = some r applyOp r1 r2 "-" = some r) : evalArith_add (ts1 ++ "+" :: ts2) r -- Noncomputable spec to evaluate an expression def do_algebra (input : List String) (result : Int) : Prop := evalArith_add input result
@@ A.4 Baseline Prompts
Snippets of the few-shot specification generator’s system and example prompts are shown in Figure 11 and Figure 12. Snippets of the few-shot isomorphism prover’s system and example prompts are shown in Figure 13 and Figure 14. COPRA’s system prompt, used for both isomorphism and correctness, is nearly identical to the original one in the COPRA paper [31]. Snippets of COPRA’s example prompt for isomorphism are shown in Figure 15.
Snippets of the few-shot implementation generator’s system and example prompts are shown in Figure 16 and Figure 17. Snippets of the few-shot correctness prover’s system and example prompts are shown in Figure 18 and Figure 19. Snippets of COPRA’s example prompt for correctness are shown in Figure 20.
{mdframed}
[roundcorner=10pt]
You are a good Lean 4 programmer. You are given a natural language specification of a function (mentioned in as a python docstring). Your task is to generate a Lean 4 proposition with a mentioned signature. The proposition takes in an implementation and program input as parameters. The proposition should hold true for all possible inputs in the domain, which means any preconditions should be mentioned in the specification to ensure that those cases are handled appropriately and hence the proposition is always valid if the implementation is correct.
The input usually follows the following format:
[NL DESCRIPTION]
def <function_name>(<input_type>) -> <output_type>
<NL Description>
Followed by the specification signature:
[SPECIFICATION SIGNATURE]
def <function_name> (impl : <function_signature>) (input : <input_type>) : Prop :=
You can first think about the problem in a general way and then write the proposition. You can also use the following template to help you with the proposition generation:
[THOUGHTS]
The proposition should be a function that takes in an implementation and input
We can use the preconditions mentioned via implication to ensure that implementations correctness
is only checked for the valid inputs ....
[END THOUGHTS]
Finally, write the generated specification in the following format:
[GENERATED SPECIFICATION]
-- Change the following lines with actual generated formal specification
(x : <input_type>), <precondition> <postcondition>
[END]
Please closely follow the format as shown in the examples below. Make sure that your response always ends with [END]. Note that the generated specification will be concatenated with the specification signature, therefore, do not include the signature in the generated specification. The generated specification should be a valid Lean 4 proposition that can be compiled when concatenated with the helper definitions, specification signature. DO NOT ever use the in keyword, it is not a valid keyword in Lean 4. Please do NOT use sorry in your proof anywhere. The proof must be complete and valid.
@@
{mdframed}
[roundcorner=10pt]
`example_user`
[NL DESCRIPTION]
def find_magnitude(x: int) -> int
Given an integer x, your task is to find the magnitude of x.
The magnitude of an integer is defined as the absolute value of the integer.
[SPECIFICATION SIGNATURE]
def generated_spec
-- function signature
(impl: Int Int)
-- inputs
(x: Int) : Prop :=
`example_assistant`
[THOUGHTS]
We need to find absolute value of an integer.
Since absolute value is always defined for all integers, we dont need to check for any preconditions.
We can write a specification which return x if x is greater than or equal to 0, otherwise -x.
It is also easy to see that program will always terminate for all integers. However, it is better to mention that in the specification.
[END THOUGHTS]
[GENERATED SPECIFICATION]
result, impl x = result
(x >= 0 result = x)
(x < 0 result = -x)
[END]
@@
{mdframed}
[roundcorner=10pt]
You are a good Lean 4 programmer. You are given:
1. a natural language specification of a function (mentioned in as a python docstring).
2. a corresponding problem specification in lean 4.
3. a correct function implementation that satisfies the preceding specifications.
Your task is to write a formal proof in Lean 4 that the function implementation is correct and satisfies the formal specification.
The correctness statement is stated in the following format:
1. First we state the natural language description of the function in a docstring format:
[NL DESCRIPTION]
def <function_name>(<input_type>) -> <output_type>
<NL Description>
4. Finally, the correctness theorem statement in Lean 4:
[THEOREM STATEMENT]
theorem correctness
(input: <input_type>)
: problem_spec implementation input
You can first think about the problem in a general way and then write the proof. You can also use the following template to help you with the proof generation:
[THOUGHTS]
The implementation should match the problem_spec in the first case because ....
The implementation should match the problem_spec in the second case because ....
[END THOUGHTS]
Finally, write a proof in Lean 4 that the implementation is correct and satisfies the specification. The proof should be in the following format:
[PROOF]
-- proof tactics
by
rw [...]
simp ...
-- more proof tactics
[END]
Please closely follow the format as shown in the examples below. Make sure that your response always ends with [END]. The proof must be in Lean 4 and should be valid. Make sure to always start your proof with ‘[PROOF]‘ followed by a by keyword. The proof should end with ‘[END]‘. DO NOT forget to write the PROOF section. The proof section must be there in your response. Please do NOT use sorry in your proof anywhere. The proof must be complete and valid.
@@ 4. Finally, the correctness theorem statement in Lean 4:
{mdframed}
[roundcorner=10pt]
`example_user`
[NL DESCRIPTION]
def find_magnitude(x: int) -> int
""" Given an integer x, your task is to find the magnitude of x.
The magnitude of an integer is defined as the absolute value of the integer. """
[GROUND TRUTH SPECIFICATION]
def problem_spec
-- function signature
(impl: Int Int)
-- inputs
(x: Int) :=
-- spec
let spec (result: Int) :=
(result 0)
(impl (-x) = result)
(result = 0 x = 0)
(0 x impl (-x) + result = 2 * x)
(x 0 impl (-x) + result = -2 * x)
-- program terminates
result, impl x = result
-- return value satisfies spec
spec result
[GENERATED SPECIFICATION]
def generated_spec
[THEOREM STATEMENT]
theorem spec_isomorphism:
impl,
(∀ x, problem_spec impl x)
(∀ x, generated_spec impl x) :=
`example_assistant`
[THOUGHTS]
To prove the equivalence of the two specifications, we can start by analyzing the generated specification.
The original specification doesnt reveal the implementation details, but we can see that the generated specification is a simple if-else statement that checks if x is less than 0.
We can use this information to write a proof that shows that the two specifications are equivalent.
The idea is to try different cases for x, and show that the properties of the generated specification hold true for the original specification as well.
[END THOUGHTS]
[PROOF]
by
unfold problem_spec
unfold generated_spec
simp
intro impl
apply Iff.intro
intro h_prob_spec
intro x
by_cases h_x_lt_0: x < 0
-- if x < 0 then
linarith
[END]

### v4

@@ 4 Related Work
Benchmarks. Recent efforts have developed benchmarks for formal verification with the onset of powerful neural models. FVAPPS [9] uses an LLM on scraped competition problems to automatically create formal specifications for 4715 problems, 1083 of which are guarded with test cases. However, the formal specifications themselves are often easily hackable (see Section˜A.1), with verification correctness guarded by a layer of test cases. Here, we aim to provide complete formal specifications, which cannot be done accurately with automatic annotation. miniCodeProps [26] contains 201 verification problems regarding data structures and induction problems; however, they do not include specification synthesis or equivalence tests. DafnyBench [27] is a benchmark of 782 stand-alone Dafny programs collected from prior benchmarks and Dafny repositories, where the synthesis task is to generate the verification conditions that allow Dafny to prove correctness. At the time, the best model was Claude 3 Opus which solved $\approx$ 68 % of the problems. Software engineering benchmarks have become extremely popular in recent literature, including benchmarking performance fixing real-world issues [14] and contamination-free code generation [12]. In our work, we employ HumanEval [4] to create Clever, our formal verification and synthesis benchmark. Formal verification is also applied in mathematical domains. Mathlib [28] and the Archive of Formal Proofs [1] constitute formal mathematical repositories in Lean and Isabelle respectively, from which benchmarks have been derived [10, 13]. ProofNet [3] serves as a benchmark for producing proper specifications of mathematical problems. PutnamBench [32] is a formal benchmark of undergraduate-level competition problems in Lean, Isabelle, and Coq.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
This work was supported by NSF awards CCF-2212559 and CCF-2403211, and a 2025 Renaissance Philanthropy AI for Math award.
@@ A.1 FVAPPS Benchmark
The FVAPPS benchmark [9] is another code generation benchmark in Lean. However, unlike Clever, which requires a comprehensive proof of full program behavior, FVAPPS only requires the proof of a limited selection of properties of the program. The limitations of this are illustrated by the FVAPPS example in Figure˜8. Here, a problem with a relatively complex natural language description only requires verifying lower-bound and upper-bound properties of the program implementation, as well as a few simple base cases. As can be seen, these properties are provably satisfied by a trivial program that always outputs 0 regardless of the input. Thus, it is clear that only requiring the proof of a small handful of properties does not capture the full intent of the natural language problem. This highlights the necessity of a verified code generation benchmark to require proofs of full program behavior, not just program properties.
/-- solve_elections: There are n voters, and two ways to convince each of them to vote for you. The first way to convince the \$i\$-th voter is to pay him \$p_i\$ coins. The second way is to make \$m_i\$ other voters vote for you, and the \$i\$-th voter will vote for free. Moreover, the process of such voting takes place in several steps. For example, if there are five voters with \$m_1 = 1\$, \$m_2 = 2\$, \$m_3 = 2\$, \$m_4 = 4\$, \$m_5 = 5\$, then you can buy the vote of the fifth voter, and eventually everyone will vote for you. Set of people voting for you will change as follows: \$5\$ \$1, 5\$ \$1, 2, 3, 5\$ \$1, 2, 3, 4, 5\$. Calculate the minimum number of coins you have to spend so that everyone votes for you. -/ def solve_elections (n : Nat) (voters : List (Nat × Nat)) : Nat := 0 theorem solve_elections_nonnegative (n : Nat) (voters : List (Nat × Nat)) : solve_elections n voters >= 0 := by rfl theorem solve_elections_upper_bound (n : Nat) (voters : List (Nat × Nat)) : solve_elections n voters <= List.foldl acc (pair : Nat × Nat) => acc + pair.2) 0 voters := Nat.zero_le _ theorem solve_elections_zero_votes (n : Nat) (voters : List (Nat × Nat)) : (List.all voters (fun pair => pair.1 = 0)) -> solve_elections n voters = 0 := fun _ => rfl theorem solve_elections_single_zero_vote : solve_elections 1 [(0, 5)] = 0 := by rfl
@@ A.2 Hard to write Specifications
Figure˜9 shows some problems for which the formal specification or the implementation is hard to write.
(a) def problem_spec -- function signature (implementation: List Rat Rat) -- inputs (xs: List Rat) := -- spec let spec (result: Rat) := let eps := (1: Rat) / 1000000; xs.length 1 xs.length % 2 = 0 poly : Polynomial Rat, poly.degree = some (xs.length - 1) (∀ i, i xs.length - 1 poly.coeff i = xs.get! i) |poly.eval result| eps; -- program termination result, implementation xs = result spec result -- possible implementation using Newtons method def implementation (xs: List Rat) : Rat := let rec poly (xs: List Rat) (x: Rat) := xs.reverse.foldl acc a => acc * x + a) 0; let rec poly (xs: List Rat) (x: Rat) := (xs.drop 1).reverse.foldl acc a => acc * x + a) 0; let rec eps := (1: Rat) / 1000000; let rec find_zero (xs: List Rat) (guess: Rat) (fuel: Nat) := let eval := poly xs guess; let eval := poly xs guess; if eval eps fuel = 0 then (guess, fuel) else let guess := (eval * guess - eval) / eval’; find_zero xs guess (fuel - 1); (find_zero xs 1.0 1000000).1 (b) def problem_spec -- function signature (implementation: Nat Nat) -- inputs (n: Nat) := -- spec let spec (result: Nat) := n > 0 (∃ i, Nat.fib i = result Nat.Prime result (∃! S : Finset Nat, S.card = n - 1 (∀ y S, (∃ k, y = Nat.fib k) y < result Nat.Prime y)) ) -- implementation without proof of -- termination def implementation (n: Nat) : Nat := let rec fib_prime (n: Nat) (i: Nat) : Nat := if Nat.Prime (Nat.fib i) then if n = 1 n = 0 then Nat.fib i else fib_prime (n - 1) (i + 1) else fib_prime n (i + 1) termination_by n decreasing_by -- Proof of termination is open problem sorry sorry fib_prime n 0
@@ A.3 Writing non-computable specifications
Figure˜10 shows a computable vs non-computable version of the specification for finding the $n^{th}$ Fibonacci number. It can be observed that the computable version of the specification leaks the implementation in contrast to the non-computable version. The non-computable specification uses an inductive definition of a recursive function.
(a) -- computable spec def problem_spec -- function signature (implementation: List Nat Nat) -- inputs (n: Nat) := -- spec let spec (result: Nat) := (n = 0 result = 0) (n = 1 result = 1) (2 n fib_array : List Nat, fib_array.length = n + 1 fib_array[0]! = 0 fib_array[1]! = 1 (∀ i, 1 < i i < n + 1 fib_array[i]! = fib_array[i - 1]! + fib_array[i - 2]!) result = fib_array[n]!) -- program termination result, implementation xs = result spec result (b) -- non-computable spec inductive fibonacci_non_computable : Prop | base0 : fibonacci_non_computable 0 0 | base1 : fibonacci_non_computable 1 1 | step : n f f₂, fibonacci_non_computable n f fibonacci_non_computable (n + 1) f fibonacci_non_computable (n + 2) (f + f₂) def problem_spec -- function signature (implementation: Nat Nat) -- inputs (n: Nat) := -- spec let spec (result: Nat) := fibonacci_non_computable n result -- program termination result, implementation xs = result spec result
@@
Writing non-computable specifications is a non-trivial task that requires a deep understanding of the problem. Figure˜11 (problem 160) presents another complex example illustrating the difficulty of formulating such specifications. Figure˜11 shows two versions of a specification for evaluating an expression given as a list of strings (["2","+","3","*","4","-","5"]). Figure˜11(a) evaluates the expression and later checks if the output matches the result (not specified in the figure), which is computable. Figure˜11(b) shows a non-computable version of the specification that checks if the result matches the output of evaluating the expression without leaking the implementation. One can notice that we need multiple inductive recursive definitions to ensure that the specification is clean and non-computable.
@@
(a) inductive Op where | add | sub | mul | floordiv deriving Repr, DecidableEq def parseOp : String Option Op | "+" => some .add | "-" => some .sub | "*" => some .mul | "//" => some .floordiv | _ => none def precedence : Op Nat | .mul | .floordiv => 2 | .add | .sub => 1 def apply : Op Int Int Int | .add, a, b => a + b | .sub, a, b => a - b | .mul, a, b => a * b | .floordiv, a, b => a / b inductive Tok where | num : Int Tok | op : Op Tok deriving Repr def tokenize : List String Option (List Tok) | [] => some [] | s :: t => match parseOp s with | some o => (tokenize t).map (Tok.op o :: ·) | none => s.toInt?.bind (fun n => (tokenize t).map (Tok.num n :: ·)) partial def evalPass (xs : List Tok) (ops : List Op) : List Tok := match xs with | Tok.num a :: Tok.op o :: Tok.num b :: rest => if o ops then evalPass (Tok.num (apply o a b) :: rest) ops else Tok.num a :: Tok.op o :: evalPass (Tok.num b :: rest) ops | x :: xs => x :: evalPass xs ops | [] => [] def evalTokens (tokens : List Tok) : Option Int := let result := $[$.mul, .floordiv], [.add, .sub]].foldl evalPass tokens match result with | [Tok.num n] => some n | _ => none def do_algebra (input : List String) : Option Int := tokenize input >>= evalTokens (b) def applyOp (x y : Int) : String Option Int | "+" => some (x + y) | "-" => some (x - y) | "*" => some (x * y) | "//" => if y == 0 then none else some (x / y) | _ => none inductive evalArith_pass : List String Int Prop | num {s : String} {n : Nat} (h : s.toNat! = n) : evalArith_pass [s] (Int.ofNat n) | binOp {ts1 ts2 : List String} {op : String} {r1 r2 r : Int} (h1 : evalArith_pass ts1 r1) (h2 : evalArith_pass ts2 r2) (hop : applyOp r1 r2 op = some r) : evalArith_pass (ts1 ++ op :: ts2) r inductive evalArith_mul : List String Int Prop | of_pass {ts r} (h : evalArith_pass ts r) : evalArith_mul ts r | step {ts1 ts2 r1 r2 r} (h1 : evalArith_mul ts1 r1) (h2 : evalArith_mul ts2 r2) (hop : applyOp r1 r2 "*" = some r applyOp r1 r2 "//" = some r) : evalArith_mul (ts1 ++ "*" :: ts2) r inductive evalArith_add : List String Int Prop | of_mul {ts r} (h : evalArith_mul ts r) : evalArith_add ts r | step {ts1 ts2 r1 r2 r} (h1 : evalArith_add ts1 r1) (h2 : evalArith_add ts2 r2) (hop : applyOp r1 r2 "+" = some r applyOp r1 r2 "-" = some r) : evalArith_add (ts1 ++ "+" :: ts2) r -- Noncomputable spec to evaluate an expression def do_algebra (input : List String) (result : Int) : Prop := evalArith_add input result
@@ A.4 Baseline Prompts
Snippets of the few-shot specification generator’s system and example prompts are shown in Figure˜12 and Figure˜13. Snippets of the few-shot isomorphism prover’s system and example prompts are shown in Figure˜14 and Figure˜15. COPRA’s system prompt, used for both isomorphism and correctness, is nearly identical to the original one in the COPRA paper [31]. Snippets of COPRA’s example prompt for isomorphism are shown in Figure˜16.
Snippets of the few-shot implementation generator’s system and example prompts are shown in Figure˜17 and Figure˜18. Snippets of the few-shot correctness prover’s system and example prompts are shown in Figure˜19 and Figure˜20. Snippets of COPRA’s example prompt for correctness are shown in Figure˜21.
You are a good Lean 4 programmer. You are given a natural language specification of a function (mentioned in as a python docstring). Your task is to generate a Lean 4 proposition with a mentioned signature. The proposition takes in an implementation and program input as parameters. The proposition should hold true for all possible inputs in the domain, which means any preconditions should be mentioned in the specification to ensure that those cases are handled appropriately and hence the proposition is always valid if the implementation is correct. The input usually follows the following format: ‘‘‘ [NL DESCRIPTION] def <function_name>(<input_type>) -> <output_type> """ <NL Description> """ ‘‘‘ Followed by the specification signature: ‘‘‘ [SPECIFICATION SIGNATURE] def <function_name> (impl : <function_signature>) (input : <input_type>) : Prop := ‘‘‘ You can first think about the problem in a general way and then write the proposition. You can also use the following template to help you with the proposition generation: ‘‘‘ [THOUGHTS] The proposition should be a function that takes in an implementation and input We can use the preconditions mentioned via implication to ensure that implementations correctness is only checked for the valid inputs .... [END THOUGHTS] ‘‘‘ Finally, write the generated specification in the following format: ‘‘‘ [GENERATED SPECIFICATION] -- Change the following lines with actual generated formal specification (x : <input_type>), <precondition> <postcondition> [END] ‘‘‘ Please closely follow the format as shown in the examples below. Make sure that your response always ends with [END]. Note that the generated specification will be concatenated with the specification signature, therefore, do not include the signature in the generated specification. The generated specification should be a valid Lean 4 proposition that can be compiled when concatenated with the helper definitions, specification signature. DO NOT ever use the in keyword, it is not a valid keyword in Lean 4. Please do NOT use sorry in your proof anywhere. The proof must be complete and valid.
@@
`example_user` [NL DESCRIPTION] def find_magnitude(x: int) -> int """ Given an integer x, your task is to find the magnitude of x. The magnitude of an integer is defined as the absolute value of the integer. """ [SPECIFICATION SIGNATURE] def generated_spec -- function signature (impl: Int Int) -- inputs (x: Int) : Prop := `example_assistant` [THOUGHTS] We need to find absolute value of an integer. Since absolute value is always defined for all integers, we dont need to check for any preconditions. We can write a specification which return x if x is greater than or equal to 0, otherwise -x. It is also easy to see that program will always terminate for all integers. However, it is better to mention that in the specification. [END THOUGHTS] [GENERATED SPECIFICATION] result, impl x = result (x >= 0 result = x) (x < 0 result = -x) [END]
@@
You are a good Lean 4 programmer. You are given: 1. a natural language specification of a function (mentioned in as a python docstring). 2. a corresponding problem specification in lean 4. 3. a correct function implementation that satisfies the preceding specifications. Your task is to write a formal proof in Lean 4 that the function implementation is correct and satisfies the formal specification. The correctness statement is stated in the following format: 1. First we state the natural language description of the function in a docstring format: ‘‘‘ [NL DESCRIPTION] def <function_name>(<input_type>) -> <output_type> """ <NL Description> """ ‘‘‘ ........ 4. Finally, the correctness theorem statement in Lean 4: ‘‘‘ [THEOREM STATEMENT] theorem correctness (input: <input_type>) : problem_spec implementation input := ‘‘‘ You can first think about the problem in a general way and then write the proof. You can also use the following template to help you with the proof generation: ‘‘‘ [THOUGHTS] The implementation should match the problem_spec in the first case because .... The implementation should match the problem_spec in the second case because .... .... [END THOUGHTS] ‘‘‘ Finally, write a proof in Lean 4 that the implementation is correct and satisfies the specification. The proof should be in the following format: ‘‘‘ [PROOF] -- proof tactics by rw [...] simp ... -- more proof tactics [END] ‘‘‘ Please closely follow the format as shown in the examples below. Make sure that your response always ends with [END]. The proof must be in Lean 4 and should be valid. Make sure to always start your proof with ‘[PROOF]‘ followed by a by keyword. The proof should end with ‘[END]‘. DO NOT forget to write the PROOF section. The proof section must be there in your response. Please do NOT use sorry in your proof anywhere. The proof must be complete and valid.
@@
`example_user` [NL DESCRIPTION] def find_magnitude(x: int) -> int """ Given an integer x, your task is to find the magnitude of x. The magnitude of an integer is defined as the absolute value of the integer. """ [GROUND TRUTH SPECIFICATION] def problem_spec -- function signature (impl: Int Int) -- inputs (x: Int) := -- spec let spec (result: Int) := (result 0) (impl (-x) = result) (result = 0 x = 0) (0 x impl (-x) + result = 2 * x) (x 0 impl (-x) + result = -2 * x) -- program terminates result, impl x = result -- return value satisfies spec spec result [GENERATED SPECIFICATION] def generated_spec ........ [THEOREM STATEMENT] theorem spec_isomorphism: impl, (∀ x, problem_spec impl x) (∀ x, generated_spec impl x) := `example_assistant` [THOUGHTS] To prove the equivalence of the two specifications, we can start by analyzing the generated specification. The original specification doesnt reveal the implementation details, but we can see that the generated specification is a simple if-else statement that checks if x is less than 0. We can use this information to write a proof that shows that the two specifications are equivalent. The idea is to try different cases for x, and show that the properties of the generated specification hold true for the original specification as well. [END THOUGHTS] [PROOF] by unfold problem_spec unfold generated_spec simp intro impl apply Iff.intro intro h_prob_spec intro x by_cases h_x_lt_0: x < 0 -- if x < 0 then ........ linarith [END]